The side length of a square can be expressed by the formula s=√A where A is the area of thw square. I f the area of the square is 50m², what is the length of its side?


1
Expert's answer
2021-03-01T07:13:50-0500

Since the area of th e square is 50m2, then

A= 50m2.

From the formula "s=\sqrt A" ,

The side length "s=\sqrt 50" m

"s=\sqrt (5 \times 5\times 2)" m

so, "s=5\sqrt 2" m.

or "s\approx 7.0710678" m